A solution of hydrochloric acid of mass fraction 18,0 % has a density of 1,0878 g/cm^3. What are the molar and mass concentration of this solution ?
1
Expert's answer
2017-11-25T09:09:07-0500
First, let's pick a 1 L sample of HCl solution with mass fraction of HCl = 18.0%. The first thing you need to do is determine is how much the 1 L sample weighs. Since density was given to you g/cm3, it'll be best to transform it in g/dm3, since 1 L = 1 dm3, so 1.0878 g/cm3*(1000 cm3/1 dm3)=1087.8 g/dm3. Mass fraction is the ratio of one substance with mass m to the mass of the total mixture, so 18%=(msolute/msolution)*100; msolute = (msolution*18)/100=(1087.8*18)/100=195.804 g HCl. There is 195.804 g*(1 mole/36.5 g)=5.3645 mol HCl. Molar concentration C =n/V=5.3645 mol / 1 L=5.3645 M (mol/L). Then mass concentration is defined as the mass of HCl divided by the volume of the mixture, so ρ=m/V=195.804 g/1 L = 195.804 g/L

Answer: 5.3645 mol/L, 195.804 g/L